. (3 pts) A small area in front of a building is triangular in shape. The perimeter of the triangle is 37 meters. The second side is one-half of the first side in length. The third side is 3 meters less than the first side. Find the length in meters of each side of the triangular region.
37= (½ * x) + (x-3)
37= 1/2x + (x-3)
39.5=2x
X=19.75
1st side= 19.75 meter
2nd side= 0.5 * 19.75 = 9.875 meters
3rd side= 19.75-3= 16.75 meters

Answer by stanbon(48535)

A small area in front of a building is triangular in shape. The perimeter of the triangle is 37 meters. The second side is one-half of the first side in length. The third side is 3 meters less than the first side. Find the length in meters of each side of the triangular region.
---------------
1st side: x meters
2nd side: x/2 meters
3rd side: x-3 meters
------------------------
EQUATION:
x + x/2 + x-3 = 37
2x + x + 2x-6 = 74
5x = 80
x = 16 meters
x/2 = 8 meters
x-2 = 13 meters
---------
Checking
16+8+13 = 37 meters
===================
Cheers,
Stan H.